A partical travels along a straight path covers 1/3 of a total time with velocity V0 and remaining 2/3 time with velocity V1 for half the distance and velocity V2 for other half of the distance Calculate its mean velocity??

Grade:11

1 Answers

m srinivas
7 years ago
For v0 velocity time taken is 1/3T
Therefore,s=voT/3
S/2v1+s/2v2=(2/3)T
S=(4v1v2/3v1+3v2)T
Mean velocity =total distance / total time
                          =(voT/3+(4v1v2/3v1+3v2)T)/T
                           =(v1v2+vov2+4v1v2)/3v1+3v2